OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

parallel line should have the same coefficients of x and y as original and so y=3x-6 and y=3x-14 are parallel to the given line but y=3x-14 is the line which is passing through the point (4,-2).So answer is y=3x-14 0r 3x-y=14

OpenStudy (anonymous):

thank you! could i ask you for one more?

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x − 10 y = x + 10 y = −x − 10 y = −x + 10